The measure of the angles of a triangle are in the extended ratio 17:16:12. What is the measure of the largest angle? A.34 B. 45 C. 48 D.68
Answer by stanbon(75887)

The measure of the angles of a triangle are in the extended ratio 17:16:12. What is the measure of the largest angle?
-----
17x + 16x + 12x = 180
45x = 180
x = 4
----
largest = 17x = 17*4 = 68 degrees
----------
Cheers,
Stan H.
--------------
A.34 B. 45 C. 48 D.68